OpenStudy (anonymous):

pleasee help!!! Finding the angle of rotation of an ellipse from its general equation ''identify the graph of the equation 2x^2+9xy+14y^2-5=0. then find θ''

OpenStudy (anonymous):

@Loser66 do u know how to do this?

OpenStudy (anonymous):

@hoblos

OpenStudy (accessdenied):

What have you tried with this question so far?

OpenStudy (anonymous):

the graph is an ellipse and i have to use this formula to find theta : Tan2θ=B / A - C soo... tan2θ= 9 / 2-14 tan2θ= -0,75

OpenStudy (accessdenied):

I believe that formula looks correct to what I recall, along with the work.. Alright, so you have: \( \displaystyle \tan 2\theta = - \frac{3}{4} \) You could find the arctangent of each side and then divide by 2. \( \displaystyle \theta = \frac{1}{2} \tan^{-1} \left( -\frac{3}{4} \right) \) This could be done on a calculator, or there might be an exact answer using trigonometry. Shall i go into that?

OpenStudy (anonymous):

hmm...how do you do tan^-1 on the calculator

OpenStudy (accessdenied):

Do you have a key for tangent? It is usually 2nd + tangent on mine.

OpenStudy (anonymous):

yeah found it! thank you SOOO much!!!! <3

OpenStudy (accessdenied):

You're welcome! :)